OpenStudy (anonymous):

verify the identity. sin(pi/2 + x)=cos x

OpenStudy (anonymous):

you can use the "addition angle formula" \[\sin(\alpha +\beta)=\sin(\alpha)\cos(\beta)+\cos(\alpha)\sin(\beta)\] with \(\alpha =\frac{\pi}{2}\) and \(\beta=x\)

OpenStudy (anonymous):

also you need to know that \(\sin(\frac{\pi}{2})=1\) and \(\cos(\frac{\pi}{2})=0\)

OpenStudy (anonymous):

what about find all solutions of the equation in the interval (0,2pi). sin(x + pi/6) - sin(x - pi/6)= 1/2